OpenStudy (anonymous):

An open box is formed from a square piece of cardboard, by removing squares of side 10 in. from each corner and folding up the sides. If the volume of the carton is then 80 in3, what was the length of a side of the original square of cardboard?

OpenStudy (anonymous):

volume is a^3. 80=a^3 a=2*cubert10. Aoriginal is 20+2*cubert10 ?

OpenStudy (anonymous):

multiple choice: a. 10+4*sqrt5 in. b. 20+4*sqrt5 in. c. 20+2*sqrt2 in. d. 10+2*sqrt2 in.